# Contact non-squeezing in lens spaces

### v1

@@ 5. Estimation of the contact capacity
Suppose that $M$ is a $\mathbb{S}^{1}$-principle bundle over a closed symplectic manifold $(W,\Omega)$, that $M$ carries a contact form $\alpha$ satisfying $\pi^{*}\Omega=\mathrm{d}\alpha$, where $\pi:M\to W=M/\mathbb{S}^{1}$ denotes the projection, and that moreover the $\alpha$-Reeb flow induces the $\mathbb{S}^{1}$-action. Then for any function $H:[0,1]\times W\to\mathbb{R}$ the $\alpha$-Hamiltonian function
@@
generates a contact isotopy $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ which is $\mathbb{S}^{1}$-equivariant, i.e. $\phi_{h}^{t}\circ\phi_{\alpha}^{s}=\phi_{\alpha}^{s}\circ\phi_{h}^{t}$ for all $s\in\mathbb{R}$, $t\in[0,1]$. Moreover $\pi\circ\phi_{h}^{t}=\psi_{H}^{t}\circ\pi$ for all $t\in[0,1]$, where $(\psi_{H}^{t})\subset\mathrm{Ham}(W,\Omega)$ is the isotopy generated by $H$.
@@ 5.1. Estimation of the lower bound
The proof of Proposition 5.2 will be an adaptation of the proof of [25, Lemma 3, Chapter 3] stating that the Hofer-Zehnder capacity of $(\mathbb{B}^{2n}(r),\omega_{0})$ is bounded from below by $\pi r^{2}$.
@@
where $J$ is the standard complex structure of $\mathbb{R}^{2n}$555after identifying $\mathbb{R}^{2n}$ with $\mathbb{C}^{n}$ via $\Phi$ defined in the introduction $J:=\Phi^{-1}\circ i\circ\Phi$ corresponds to the multiplication by $i$ and $\nabla H$ denotes the gradient of $H$ with respect to the standard Euclidean scalar product. A direct computation shows that for all $t\in[0,1]$ and $x\in\mathbb{B}^{2n}(r)$
@@
The relation (9) actually holds for all $T\in[0,1]$ since by [26, Thm 1] the radius $r\leq\sqrt{2}$ and so $\delta=1$ (see also Remark 1.2). The $\alpha_{k}$-Hamiltonian function $h$ being supported in $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n},(r)))$ we get by definition that
@@ 5.2. Estimation of the upper bound
Let us first consider $(\mathbb{R}^{2}=\mathbb{R}^{2}\times\{0\},dx_{1}\wedge dy_{1})$ the symplectic submanifold of $(\mathbb{R}^{2n},\sum\limits_{i=1}^{n}dx_{i}\wedge dy_{i})$. We recall a Lemma from [24, Lemma 4.2].
@@
where $\widetilde{\phi}:=[(\phi_{t})]\in\widetilde{\mathrm{Cont}_{0}}(L_{K}^{2n+1},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $L_{k}^{2n+1}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, the last point of Proposition 2.3 implies that
@@
6. Possible generalization to orderable prequantizations
6.1. Orderability and pseudo-spectral selectors
Let $(M,\xi)$ be a closed coorientable contact manifold and from now on we fix a coorientation of $\xi$, i.e. a vector field $X\in\chi(M)$ so that $X(M)\subset TM\setminus\xi$. We say that a contact form $\alpha$ supports $\xi$ if $\ker\alpha=\xi$ and $\alpha(X)>0$. A contact isotopy $(\phi_{t})$ is said to be non-negative if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha(\frac{d}{dt}\phi_{t}(x))\geq 0$ for all $(t,x)\in[0,1]\times M$. Recall from [15] that $(M,\xi)$ is said to be orderable if the only non-negative contact isotopy $(\phi_{t})$ representing $\mathrm{id}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is the constant isotopy. Under the assumption that $(M,\xi)$ is orderable, in [3, 5, 30] the authors show that for any contact form $\alpha$ supporting $\xi$ the map $c_{+}^{\alpha}:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}$ defined by
[…]
is a well defined $C^{1}$-continuous map, where $\widetilde{\varphi}\preceq\widetilde{\psi}$ if and only if $\widetilde{\psi}\cdot\widetilde{\varphi}^{-1}$ can be represented by a non-negative contact isotopy $(\phi_{t})$. Moreover it is shown in [3, 30] that this map shares a lot of properties with the $\alpha$-spectral selector of Proposition 2.3 constructed in [4]. Namely for all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$
@@ 6.1. Orderability and pseudo-spectral selectors
    $\widetilde{\varphi}\preceq\widetilde{\psi}$ implies that $c^{\alpha}_{+}(\widetilde{\varphi})\leq c^{\alpha}_{+}(\widetilde{\psi})$
@@
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $c^{\alpha}_{+}(\widetilde{\varphi})>0$
@@
    $c_{+}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{+}^{\alpha}(\widetilde{\psi})+c_{+}^{\alpha}(\widetilde{\varphi})$
@@
    $\int_{0}^{1}\min\alpha\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t\leq c_{+}^{\alpha}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t.$
As in Section 3 defining a map $C_{\alpha}:\{\text{non-empty open set of }M\}\to\mathbb{R}_{>0}\cup\{+\infty\}$
$U\mapsto C_{\alpha}(U):=\sup\{c_{+}^{\alpha}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi)\},$
we get that
if $\text{if }U\subset V\text{ then }C_{\alpha}(U)\leq C_{\alpha}(V).$ (10)
Without making more assumption on the contact manifold $(M,\xi)$ and the map $c_{+}^{\alpha}$ one cannot say much more about $C_{\alpha}$. However in [3] the authors made the following conjecture
Conjecture 6.1.
If $(M,\xi)$ is a closed orderable contact manifold then $c_{+}^{\alpha}$ is an $\alpha$-spectral selector for any contact form $\alpha$ supporting $\xi$.
The next Lemma gives an important feature of $C_{\alpha}$ under the assumptions that the conjecture 6.1 holds and that the $\alpha$-Reeb flow is periodic.

### v2

@@ 6. Estimation of the contact capacity
Let $T>0$ be any positive real number and $\pi$ be the projection of a $S^{1}:=\mathbb{R}/T$ principal bundle $M$ over a symplectic manifold $(W,\Omega)$ so that $\pi^{*}\Omega=\mathrm{d}\alpha$ for some contact form $\alpha$ on $M$, whose Reeb flow induces the $S^{1}$-action. Then for any function $H:[0,1]\times W\to\mathbb{R}$ the $\alpha$-Hamiltonian function
@@
generates a contact isotopy $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ which is $S^{1}$-equivariant, i.e. $\phi_{h}^{t}\circ\phi_{\alpha}^{s}=\phi_{\alpha}^{s}\circ\phi_{h}^{t}$ for all $s\in\mathbb{R}$, $t\in[0,1]$. Moreover $\pi\circ\phi_{h}^{t}=\psi_{H}^{t}\circ\pi$ for all $t\in[0,1]$, where $(\psi_{H}^{t})\subset\mathrm{Ham}(W,\Omega)$ is the isotopy generated by $H$.
@@ 6.1. Estimation of the lower bound
The proof of Proposition 6.2 will be an adaptation of the proof of [26, Lemma 3, Chapter 3] stating that the Hofer-Zehnder capacity of $(\mathbb{B}^{2n}(r),\omega_{0})$ is bounded from below by $\pi r^{2}$.
@@
where $J$ is the standard complex structure of $\mathbb{R}^{2n}$777after identifying $\mathbb{R}^{2n}$ with $\mathbb{C}^{n}$ via $\Phi$ defined in the introduction $J:=\Phi^{-1}\circ i\circ\Phi$ corresponds to the multiplication by $i$ and $\nabla H$ denotes the gradient of $H$ with respect to the standard Euclidean scalar product. A direct computation shows that for all $t\in[0,1]$ and $x\in\mathbb{B}^{2n}(r)$
@@
The relation (9) actually holds for all $T\in[0,1]$ since by [27, Thm 1] the radius $r\leq\sqrt{2}$ and so $\delta=1$ (see also Remark 1.2). The $\alpha_{k}$-Hamiltonian function $h$ being supported in $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n},(r)))$ we get by definition that
@@ 6.2. Estimation of the upper bound
Let us first consider $(\mathbb{R}^{2}=\mathbb{R}^{2}\times\{0\},dx_{1}\wedge dy_{1})$ the symplectic submanifold of $(\mathbb{R}^{2n},\sum\limits_{i=1}^{n}dx_{i}\wedge dy_{i})$. We recall a Lemma from [25, Lemma 4.2].
@@
where $\widetilde{\phi}:=[(\phi_{h}^{t})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $L_{k}^{2n+1}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{h}^{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, the last point of Proposition 3.2 implies that
@@
7. Order spectral selectors for strongly orderable prequantizations
7.1. Order spectral selectors on Legendrian isotopy class
Let $(M,\xi)$ be a cooriented contact manifold, not necessarily closed. As in the closed case in Section 2, a contact isotopy, that we denote by $(\phi_{t})$, is a $[0,1]$-family of contactomorphisms $\{\phi_{t}\}\subset\mathrm{Cont}(M,\xi)$ so that $[0,1]\times M\to M$ is smooth, and we denote by $\mathrm{Cont}_{0}(M,\xi)$ the set of contactomorphisms isotopic to the identity, and by $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ the ones that are isotopic to the identity through a compactly supported isotopy. Similarly, a $[0,1]$-family of closed Legendrians $(\Lambda_{t})\subset(M,\xi)$ is a Legendrian isotopy if and only if there exists a compactly supported contact isotopy $(\phi_{t})\subset\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ parametrizing it, i.e. $\phi_{t}(\Lambda_{0})=\Lambda_{t}$ and $\phi_{t}$ is compactly supported. Note that all the Legendrians we deal with in this paper are closed (even if the contact manifold itself may not). Therefore the isotopy class of a closed Legendrian $\Lambda\subset M$, that we denote by $\mathcal{L}(\Lambda)$, is an homogeneous space $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)/\mathrm{Stab}(\Lambda)$ and inherits the $C^{1}$-topology of $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$. We denote by ~$\widetilde{\mathcal{L}}(\Lambda)$ its universal cover and $\Pi:\widetilde{\mathcal{L}}(\Lambda)\to\mathcal{L}(\Lambda)$ the natural projection. Note that we have a natural left action of $\mathrm{Cont}_{0}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$888the set of equivalence classes of contact isotopy starting at the identity, where two such contact isotopies are identified if they are homotopic relatively to endpoints) on $\mathcal{L}(\Lambda)$ (resp. ~$\widetilde{\mathcal{L}}(\Lambda)$).
If $\alpha$ is a contact form whose Reeb vector field is complete, then as for contactomorphisms we have a notion of spectrum. More precisely let $\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0}\in\widetilde{\mathcal{L}}(\Lambda)$ and denote by $\Lambda_{1}$ and $\Lambda_{0}$ their respective projections to $\mathcal{L}(\Lambda)$ then
$\mathrm{Spec}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})=\{t\in\mathbb{R}\ |\ \phi_{\alpha}^{t}(\Lambda_{0})\cap\Lambda_{1}\neq\emptyset\}.$
A Legendrian isotopy $(\Lambda_{t})$ is said to be non-negative (resp. positive) if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha(\frac{d}{dt}\phi_{t}(x))\geq 0$ (resp. $>0$) for all $x\in\Lambda_{0}$ and any compactly supported contact isotopy $(\phi_{t})$ parametrizing it. We say, following [16], that ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable if there does not exist any positive contractible loop of Legendrians in $\mathcal{L}(\Lambda)$. Assuming that ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable implies that the binary relation $\preceq$, defined by $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ if and only if there exists a path $(\widetilde{\Lambda}_{t})$ joining $\widetilde{\Lambda}_{0}$ to $\widetilde{\Lambda}_{1}$ such that $(\Pi(\widetilde{\Lambda}_{t}))$ is a non-negative Legendrian isotopy, is a partial order.
For any contact form $\alpha$ supporting $\xi$ whose Reeb vector field is complete in [3] the authors defined a map $\ell_{+}^{\alpha}:\widetilde{\mathcal{L}}(\Lambda)\times\widetilde{\mathcal{L}}(\Lambda)\to\mathbb{R}\cup\{-\infty\}$
$(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\mapsto\inf\left\{t\in\mathbb{R}\ |\widetilde{\Lambda}_{1}\preceq\widetilde{\phi_{\alpha}^{t}}\cdot\widetilde{\Lambda}_{0}\right\}=\inf\left\{\underset{[0,1]\times\Lambda_{0}}{\max}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\ |\ [(\phi_{t})]\cdot\widetilde{\Lambda}_{0}=\widetilde{\Lambda}_{1}\right\},$
where $\Lambda_{0}:=\Pi(\widetilde{\Lambda}_{0})$, and showed the following:
Proposition 7.1 ([3]).
The map $\ell_{+}^{\alpha}$ takes value in $\mathbb{R}$ if and only if ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable. Moreover if ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable then for any $\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{2}\in\widetilde{\mathcal{L}}(\Lambda)$ and any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$
  1. (1)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{0})=0$
  2. (2)
    $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ implies that $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\geq 0$ and $\ell_{+}^{\alpha}(\Lambda_{0},\Lambda_{1})\leq 0$
  3. (3)
    $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ and $\widetilde{\Lambda}_{0}\neq\widetilde{\Lambda}_{1}$ implies that $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})>0$
  4. (4)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{2},\widetilde{\Lambda}_{0})\leq\ell_{+}^{\alpha}(\widetilde{\Lambda}_{2},\widetilde{\Lambda}_{1})+\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$
  5. (5)
    $\ell_{+}^{\alpha}(\widetilde{\varphi}\cdot\widetilde{\Lambda}_{1},\widetilde{\varphi}\cdot\widetilde{\Lambda}_{0})=\ell_{+}^{\varphi^{*}\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$ where $\varphi=\Pi(\widetilde{\varphi})$
  6. (6)
    For any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\underset{x\in\Pi(\widetilde{\Lambda}_{0})}{\min}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t\leq\ell_{+}^{\alpha}(\widetilde{\varphi}\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{0})\leq\int_{0}^{1}\underset{x\in\Pi(\widetilde{\Lambda}_{0})}{\max}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t$
  7. (7)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\in\mathrm{Spec}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$.
7.2. Spectral selectors for strongly orderable closed contact manifold
Let $(M,\xi=\ker\alpha)$ be a closed cooriented contact manifold and consider its contact product $(M\times M\times\mathbb{R},\Xi=\ker(\beta:=\mathrm{pr}_{2}^{*}\alpha-e^{\theta}\mathrm{pr}_{1}^{*}\alpha))$ as in Section 1.2. To any $[(\varphi_{t})]=\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ one can associate $[(\Phi_{t})]=\widetilde{\Phi}\in\widetilde{\mathrm{Cont}_{0}}(M\times M\times\mathbb{R},\Xi)$ defined by
$\Phi_{t}:(x_{1},x_{2},\theta)\mapsto(x_{1},\varphi_{t}(x_{2}),\theta+g_{t}(x_{2})),$ (10)
where $g_{t}$ is the conformal factor of $\varphi_{t}$ with respect to $\alpha$, i.e. $\phi_{t}^{*}\alpha=e^{g_{t}}\alpha$. We say that $\widetilde{\Phi}$ is the graph of $\widetilde{\varphi}$. Direct computations show that
$\mathrm{Spec}^{\alpha}(\widetilde{\varphi})=\mathrm{Spec}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta),$ (11)
where we have identified the Legendrian $\Delta=\{(x,x,0)\ |\ x\in M\}\subset(M\times M\times\mathbb{R},\ker\beta)$ with the equivalence class of the constant isotopy at $\Delta$ in ~$\widetilde{\mathcal{L}}(\Delta)$, and that for all $t\in[0,1]$ and $(x_{1},x_{2},\theta)\in M\times M\times\mathbb{R}$
$\beta\left(\frac{d}{dt}\Phi_{t}(x_{1},x_{2},\theta)\right)=\alpha\left(\frac{d}{dt}\phi_{t}(x_{2})\right).$ (12)
Suppose that in addition to be closed $(M,\xi=\ker\alpha)$ is also strongly orderable, i.e. ~$\widetilde{\mathcal{L}}(\Delta)$ is orderable, then using Proposition 7.1 we can define the map
$C_{+}^{\alpha}:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}\quad\quad\widetilde{\varphi}\mapsto\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta),\text{ where }\widetilde{\Phi}\text{ is the graph of }\widetilde{\varphi}$
which satisfies moreover the following:
Proposition 7.2.
Suppose that $(M,\xi=\ker\alpha)$ is a closed strongly orderable contact manifold. Then for any $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ denoting by $\widetilde{\Phi},\widetilde{\Psi}\in\widetilde{\mathrm{Cont}_{0}}(M\times M\times\mathbb{R},\ker\beta)$ their corresponding graphs
@@ 7.2. Spectral selectors for strongly orderable closed contact manifold
    $\widetilde{\psi}\preceq\widetilde{\varphi}$ implies that $C_{+}^{\alpha}(\widetilde{\psi})\leq C_{+}^{\alpha}(\widetilde{\varphi})$
@@
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $C_{+}^{\alpha}(\widetilde{\varphi})>0$
@@
    $C_{+}^{\alpha}(\widetilde{\psi}\widetilde{\varphi})\leq\ell_{+}^{\Psi^{*}\beta}(\widetilde{\Phi}\Delta,\Delta)+C_{+}^{\alpha}(\widetilde{\psi})$ where $\Psi:=\Pi(\widetilde{\Psi})\in\mathrm{Cont}_{0}(M\times M\times\mathbb{R},\Xi)$
@@
    $\int_{0}^{1}\min\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t\leq C_{+}^{\alpha}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t$
  6. (6)
    $C_{+}^{\alpha}(\widetilde{\varphi})\in\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$.
Proof.
  1. (1)
    $C_{+}^{\alpha}(\mathrm{id})=\ell_{+}^{\beta}(\Delta,\Delta)=0$
  2. (2)
    By the triangle inequality of Proposition 7.1 we have
    $\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\Delta)$ $\geq\ell_{+}^{\beta}(\widetilde{\Phi}(\Delta),\Delta)-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta)-\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)$
    $=-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta).$
    If we suppose that $\widetilde{\psi}\preceq\widetilde{\varphi}$ this implies, using (12), that $\widetilde{\Psi}\cdot\Lambda\preceq\widetilde{\Phi}\cdot\Lambda$ and thus using the monotonicity property of $\ell_{+}^{\beta}$ we get the desired result
    $C_{+}^{\alpha}(\widetilde{\varphi})-C_{+}^{\alpha}(\widetilde{\psi})\geq-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta)\geq 0.$
  3. (3)
    Let $\widetilde{\varphi}$ be such that $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$. Let $(\phi_{t})$ be a non-negative contact isotopy so that $[(\phi_{t})]=\widetilde{\varphi}$ and consider its graph $(\Phi_{t})$ which is also non-negative contact isotopy by (12). Since $\widetilde{\varphi}\neq\mathrm{id}$ there exists $T\in[0,1]$ and $x\in M$ so that $\phi_{T}(x)\neq x$ which implies that $\Phi_{T}(\Delta)\neq\Delta$. Therefore using the non-degeneracy and monotonicity of $\ell_{+}^{\beta}$ we get that
    $C_{+}^{\alpha}(\widetilde{\varphi})=\ell_{+}^{\beta}(\widetilde{\Phi}_{1}\cdot\Delta,\Delta)\geq\ell_{+}^{\beta}(\widetilde{\Phi}_{T}\cdot\Delta,\Delta)>0$
  4. (4)
    By the triangle inequality property and the naturality of $\ell_{+}^{\beta}$ we get that
    $\ell_{+}^{\beta}(\widetilde{\Psi}\widetilde{\Phi}\cdot\Delta,\Delta)\leq\ell_{+}^{\beta}(\widetilde{\Psi}\widetilde{\Phi}\cdot\Delta,\widetilde{\Psi}\cdot\Delta)+\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\Delta)=\ell_{+}^{\Psi^{*}\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)+C_{+}^{\alpha}(\widetilde{\psi}).$
  5. (5)
    The estimates come from the one of Proposition 7.1 combined with (12).
  6. (6)
    The spectrality of $C_{+}^{\alpha}$ follows from (11) and the spectrality of $\ell_{+}^{\beta}$.
As a corollary we get the following.
Corollary 7.3.
Suppose that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$ can be represented by a path $(\psi_{t})$ so that $\psi_{t}^{*}\alpha=\alpha$ for all $t\in[0,1]$ then
$C_{+}^{\alpha}(\widetilde{\psi}\widetilde{\varphi})\leq C_{+}^{\alpha}(\widetilde{\psi})+C_{+}^{\alpha}(\widetilde{\varphi}).$
Proof.
Fist note that for any $\psi\in\mathrm{Cont}_{0}(M,\ker\alpha)$ if $g:M\to\mathbb{R}$ is the conformal factor of $\psi$ then the conformal factor of its graph $\Psi\in\mathrm{Cont}_{0}(M\times M\times\mathbb{R},\ker\beta)$ is given by $G:M\times M\times\mathbb{R}\to\mathbb{R},\ (x_{1},x_{2},\theta)\mapsto g(x_{2})$. Therefore if $\psi^{*}\alpha=\alpha$ it implies that $\Psi^{*}\beta=\beta$. The proof of the corollary follows now directly from the triangle inequality of Proposition 7.2. ∎
Remark 7.4.
If $\psi\in\mathrm{Cont}_{0}(M,\xi)$ is such that $\psi^{*}\alpha=\alpha$ then $\psi\circ\phi_{\alpha}^{t}=\phi_{\alpha}^{t}\circ\psi$ for all $t\in\mathbb{R}$. Indeed, it implies that $\psi^{*}R_{\alpha}=R_{\alpha}$
7.3. Order spectral selectors for strongly orderable prequantizations
Let $(M,\ker\alpha)$ be a cooriented contact manifold and $\Gamma$ a finite group acting freely on $M$ by $\alpha$-strict contactomorphism, i.e. $\gamma^{*}\alpha=\alpha$ for any $\gamma\in G$. Then $\alpha$ descends to a contact form $\alpha^{\prime}$ on $M^{\prime}:=M/\Gamma$ and the subgroup $\mathrm{Cont}_{0}^{\Gamma}(M,\ker\alpha)\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ of contactomorphisms isotopic to the identity that are equivarient with respect to the action of $\Gamma$ is a covering space over $\mathrm{Cont}_{0}(M^{\prime},\ker\alpha^{\prime})$. In particular any contact isotopy starting at the identity in $\mathrm{Cont}_{0}(M^{\prime},\ker\alpha^{\prime})$ admits a unique lift to $\mathrm{Cont}_{0}^{\Gamma}(M,\ker\alpha)$ starting at the identity and so we deduce a lifting map
$\widetilde{\mathrm{Cont}_{0}}(M^{\prime},\ker\alpha^{\prime})\to\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha).$ (13)
Recall that a contact manifold $(M_{1},\xi_{1}^{\prime})$ is a prequantization over a symplectic manifold $(W,\omega^{\prime})$ if there exists a $\mathbb{T}^{1}$-principal bundle $\pi_{1}^{\prime}:M_{1}\to W$, where $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and a contact form $\alpha_{1}^{\prime}$ supporting $\xi_{1}^{\prime}$ so that the Reeb flow of $\alpha_{1}^{\prime}$ is Zoll of minimal period $2\pi$, induces the $\mathbb{T}^{1}$-action and satisfies $\pi_{1}^{\prime*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$.
From now on until the end of the paper, we drop all the primes from our notations, and $(M_{1},\xi_{1}=\ker\alpha_{1})$ is any closed prequantization over some symplectic manifold $(W,\omega)$ and $\alpha_{1}$ is the contact form whose Reeb flow is Zoll of minimal period $2\pi$, induces the $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$ action on $M_{1}$ and satisfies $\pi_{1}^{*}\omega=\alpha_{1}$, where $\pi_{1}:M_{1}\to W=M_{1}/\mathbb{T}^{1}$ is the natural projection. For any $k\in\mathbb{N}_{>0}$, $\mathbb{Z}_{k}$ is a finite group acting on $(M_{1},\xi_{1})$, by $[j]\cdot x=\phi_{\alpha_{1}}^{\frac{2\pi j}{k}}(x)$ for all $j\in\mathbb{Z}_{k},\ x\in M_{1}$, and we denote by $M_{k}:=M_{1}/\mathbb{Z}_{k}$ the quotient manifold, by $\xi_{k}$ the contact distribution given by the kernel of the induced contact $\alpha_{k}$ and by $\pi_{k}:M_{k}\to W$ the natural projection (see 1.2). We have the following Lemma.